OpenStudy (anonymous):

Solve the equation: (3x - 2)(7x + 1) = 0

OpenStudy (anonymous):

sorry, x=2/3 or x=-1/7

OpenStudy (anonymous):

3x=2----> 2/3 7x=-1-----> -1/7

OpenStudy (anonymous):

3x-2=0 or 7x+1=0

OpenStudy (anonymous):

To solve a problem like this, you set each factor to zero and solve. You roots (answers) will be the results. However, like everything, make sure you check your answers by plugging into the original problem, because sometimes the results will not be valid.

OpenStudy (callisto):

when the equation exist either (3x - 2) or (7x + 1) = 0 so (3x - 2) = 0 or (7x + 1) = 0 x = 2/3 or x=-1/7

OpenStudy (callisto):

But make sure the things on the left are factorised before you put it = 0 respectively